how long will it take a 60kg man to climb a height of 22m if he expended energy at the rate of 0.25kw? [g=ms]
1
Expert's answer
2018-03-13T16:40:08-0400
We can find the time that man needs to climb a height of 22 m from the definition of the power:
P=Fv=Fh/t=mgh/t,
t=mgh/P=(60 kg∙10 m/s^2 ∙22 m)/(0.25∙〖10〗^3 W)=52.8 s.